OpenStudy (anonymous):

4^(x+7)=6^(x+1) answers are x ≈ 0.04 x ≈ 0.05 x ≈ 19.51 x ≈ 28.35

OpenStudy (abb0t):

take the natural log. Remember that: \[\ln(a)^z=z \ln(a)\]

OpenStudy (anonymous):

so do i tske the natural numbers of 4 and 6?

OpenStudy (abb0t):

Yes.

OpenStudy (anonymous):

4s natural number is 1.38629436112 6s natural number is 1.79175946923 now what do i do?

OpenStudy (zehanz):

Take log (or ln, as @abb0t said) on both sides:\[\log 4^{x+7}=\log 6^{x+1}\]Then:\[(x+7) \log 4 = (x+1) \log 6\]Work out the brackets:\[x \log 4 + 7 \log 4 = x \log 6 + \log 6\]Collect terms:\[x(\log 4 - \log 6)=\log 6 - 7 \log 4\]Now \[x = \frac{ \log 6 - 7 \log 4 }{ \log 4 - \log 6 }\]Pull out your calculator and you'll see what the right answer is!

OpenStudy (anonymous):

i get -7 when i do that

OpenStudy (zehanz):

Be careful with the calculator. If you don't put brackets in the right places, the answer will make no sense. I typed: (log(6)-7log(4))/(log(4)-log(6)) and got 19.5141355.